1254 6B use Cramers rule to solve linear system [ 3x+2y+43 = 17... (1) 2x+4y-32=8 (2) 44. By+22=-1 (3)
Answer #1

3x+29+42 - 17 21 + 49 - 37 = 8 4-0 - 39 + २ = | 3 २ २ 4 38-9)२ (4++2) +4(-6-16) -3 4 -3 %3D -3 - 3२ - 88 = -123 Dx = 17 4 PA